OpenStudy (anonymous):

Find a rational number and an irrational number that is between 7.7 and 7.9. Include the decimal approxdimation of the irrational number to the nearest hundredth.

OpenStudy (welshfella):

the rational number could be 7.8

OpenStudy (anonymous):

I cant figure out the irrational number

OpenStudy (welshfella):

7.7^2 = 59.29 7.9^2 = 62.41 an irrational number would be square root of a number between these 2 - say sqrt 61

OpenStudy (welshfella):

= 7.81 to nearest 1/100

OpenStudy (freckles):

well pi is approx 3.14 what can you add to 3.1 to get 7.7 well 7.7-3.1=4.6 so another way to pick a irrational number between the given numbers is to say pi+4.6

OpenStudy (solomonzelman):

I saw this same question today, and I proposed the exact same exampoles of rat. and irat. numbers:D

OpenStudy (solomonzelman):

welshfella you shouldn't cheat of me;O

OpenStudy (welshfella):

pure coincidence - honest

OpenStudy (welshfella):

you know the saying - Great minds think alike!

OpenStudy (solomonzelman):

:D